OpenStudy (anonymous):

IS THIS CORRECT? |3t+6|=9 3t+6=9 3t-6=9 -6 -6 -6 -6 3t=3 3t=3 divide by 3 divide by 3 t=1 t=1 1,1 or just 1 OR would it just be the first example i can not remember if i am suppose to add and minus.

OpenStudy (anonymous):

nope this is not correct

OpenStudy (superdavesuper):

Sorry it is not....can u go back 2 the old post where i wrote out the two eqns 4 u?

OpenStudy (phi):

the | | give you 2 equations. one equation by just "dropping" the | | 3t+6=9 which you solved. t=1 the other equation is by replacing the | | with parens, and multiplying the whole thing by -1: -(3t+6)=9 You just flipped the sign on the 6. (so that answer will be wrong) solve -(3t+6)=9 or -3t - 6 = 9

OpenStudy (the_fizicx99):

I thought you would make the 9 negative to get another solution

OpenStudy (phi):

you could do that: -(3t+6)=9 multiply both sides by -1 3t+6= -9 which is another way to write the equation.

OpenStudy (the_fizicx99):

:P its much faster if you prefer :D